OpenStudy (anonymous):

The ratio of two numbers is 5 to 3. The sum of the numbers is 16. What are the numbers? 8 and 8 9 and 7 15 and 9 10 and 6

Vocaloid (vocaloid):

let x and y be the two numbers ratio of x to y = x/y = 5/3 sum of the numbers: x + y = 16 do you know what do do next?

OpenStudy (anonymous):

is it the first 1

Vocaloid (vocaloid):

nope, remember that the ratio is 5 to 3, which means that the first number divided by the second number must be 5/3

OpenStudy (anonymous):

the third?

Vocaloid (vocaloid):

no, please don't try to guess, I would like you solve the problem mathematically

Vocaloid (vocaloid):

no, here's a hint: 8/8 = ? 9/7 = ? 15/9 = ? 10/6 = ? make sure to simplify fractions

OpenStudy (anonymous):

10 & 6

OpenStudy (skullpatrol):

10/6 = (2•5)/(2•3)